# Protecting Bystander Privacy via Selective Hearing in LALMs

### v1

@@ Abstract
Large audio language models (LALMs) are increasingly deployed in real-world settings where they inevitably capture speech from unintended nearby bystanders, raising privacy risks that existing benchmarks and defences largely overlook. We introduce SH-Bench, the first benchmark designed to evaluate selective hearing: a model’s ability to attend to an intended main speaker while refusing to process or reveal information about incidental bystander speech. SH-Bench contains 3,968 multi-speaker audio mixtures spanning both real-world and synthetic scenarios, paired with 77k multiple-choice questions that probe models under general and selective operating modes. We propose Selective Efficacy (SE), a unified metric capturing both multi-speaker comprehension and bystander-privacy protection. Our evaluation of state-of-the-art open-source and proprietary LALMs reveals substantial privacy leakage, with strong audio understanding failing to translate into selective protection of bystander privacy. To mitigate this gap, we introduce Bystander Privacy Fine-Tuning (BPFT), a training pipeline that teaches models to refuse bystander-related queries without degrading main-speaker comprehension. BPFT yields substantial gains which improve SE by up to 15.9% over Gemini 2.5 Pro, demonstrating that selective hearing is learnable but far from achieved in current LALMs. SH-Bench and BPFT provide the first systematic framework for measuring and improving bystander privacy in audio foundation models. [^1] https://huggingface.co/datasets/BrianatCambridge/SelectiveHearingBench
Protecting Bystander Privacy via Selective Hearing in LALMs
@@ 1 Introduction
Large audio-language models (LALMs), especially the recent efforts including Speech-LLaMA (wu2023decoder), SALMONN (tang2023salmonn), BLSP (wang2023blsp), and Qwen-Audio (chu2023qwen), extend the capabilities of text-based large language models (LLMs) to the acoustic domain. As LALMs are deployed in real-world settings such as voice assistants and wearable devices metaglasses; geminilive, they passively capture open-domain speech in uncontrolled environments, which inevitably introduces significant privacy risks. Human voices contain sensitive acoustic attributes such as timbre, pitch, and prosody that can reveal identity, emotional state, and health conditions (nautsch2019gdpr; backstrom2023privacy; wang2025man; aloufi2021configurable). When trained on large-scale real-world speech corpora, LALMs often inadvertently memorise this information hartmann2023sok; mccoy2023much, leading to potential exposure or re-identification chen2023slmia. Moreover, prior work further shows that LALMs are vulnerable to various privacy attacks  (tseng2021membership; carlini2021extracting; yang2025can; audiojailblogs) which amplify the risks of sensitive information leakage.
@@
Figure 1: An illustration of the bystander privacy challenge in LALMs. The primary speaker interacts with a LALM, while nearby bystanders may be unintentionally recorded and could unknowingly reveal private information. To protect bystander privacy, the LALM should attend only to the primary speaker and refuse to answer any queries concerning bystanders.
@@
Yet, existing mitigation efforts focus primarily on active users (tran2023privacy; koudounas2025alexa; cheng2025speech; alexos2025defending), those who knowingly interact with the model. In contrast, a significant and overlooked group in real-world deployment contexts are bystanders: individuals whose speech is incidentally captured without their knowledge, consent, or intent to engage [^2] In social science, bystander typically refers to an observer of an event without active participation (e.g, “bystander effect” in zapata2024bystanders). Here we use the definition and research focus established in smart-device contexts yao2019privacy; saqib2025bystander, viewed through a privacy lens that highlights risks when data is passively captured or shared.. Bystanders face the same privacy risks as active users, but they neither control nor even know how their speech is processed. This disconnect raises a critical question: Can LALMs be designed to selectively attend to intended input while refusing to expose bystander information? As illustrated in Fig.1, the LALM should refuse any request targeting a bystander who may unknowingly disclose private speech.
@@
To address the critical gap that bystander privacy has not been quantified, compared, or evaluated at the behavioural level of LALMs, this paper proposes SH-Bench, a Selective Hearing Benchmark, the first benchmark for assessing the capability of LALMs to protect bystander privacy through selective hearing. SH-Bench consists of multi-speaker audio samples with five-way classification tasks where one of the option designed to be “I don’t know”. An evaluation framework is designed for SH-Bench which allows the assessment of both model comprehension abilities in multi-speaker scenarios and bystander privacy protection, with a unified criterion, the selective efficacy (SE) proposed. Moreover, we introduce a bystander-privacy fine-tuning pipeline, providing training data intended to enhance bystander protection in LALMs. Overall, our results reveal a substantial lack of bystander privacy protection in existing LALMs without fine-tuning. Main contributions of this paper are summarised as follows.
@@
  • We propose the first selective hearing benchmark, SH-Bench, for LALMs, particularly focusing on privacy protection in these LLMs of bystanders in a multi-speaker environment.
@@
  • We propose the evaluation framework for SH-Bench featuring two different operation modes and two speaking parties. We also propose SE as a unified metric balancing model comprehension abilities and bystander privacy protection.
@@
  • We propose a bystander privacy fine-tuning (BPFT) pipeline from training data curation to supervised fine-tuning to enhance bystander privacy in LALMs. BPFT achieved consistent and large improvements on bystander privacy protection, with 15.9% absolute SE improvements compared to Gemini 2.5 Pro.
@@ 2.1 Multi-Speaker Benchmarks
It is common for speech benchmarks to be either fully multi-speaker or to include substantial multi-speaker segments, as speaker diarization is one of the core representative tasks in speech processing. We categorise existing multi-speaker benchmarks into two groups: (i). Benchmarks are primarily designed for non-privacy related tasks such as automatic speech recognition (ASR), spoken question answering, speaker diarisation and speech separation. Traditional corpora in this category include AMI kraaij2005ami, LibriMix cosentino2020librimix, Switchboard godfrey1992switchboard, and the CHiME series zeinali2018deepmine; garcia2019speaker. More recent benchmarks tailored for LALMs include Dynamic-SUPERB huang2024dynamic, MMAU sakshi2024mmau, MMMU yue2024mmmu, and AudioSet-R sun2025audioset. Audio-visual multi-speaker benchmarks such as AVUT yang2025audio and AV-SUPERB tseng2024av have also been developed to evaluate tasks that require joint audio-visual understanding. (ii). Safety- and privacy-oriented audio benchmarks focus on evaluating model robustness, safety risks, and privacy leakage in multi-speaker settings. For instance, the multi-speaker anonymisation benchmark in miao2025benchmark examines privacy risks and potential mitigation strategies in overlapping conversational audio, and SACRED-Bench yang2025speech, the first multi-speaker jailbreak benchmark, contains dialogues in which harmful instructions are embedded beneath or alongside benign speech. However, these benchmarks focus solely on active speakers and overlook bystanders, leaving a critical gap that we address in this study.
2.2 Privacy Risks with LALMs
Privacy research on LALMs has largely focused on protecting the primary speaker, without distinguishing between intended users and bystanders. Early work reveals that both end-to-end ASR and self-supervised speech encoders can leak training membership information through black-box queries, demonstrating that speech representations themselves carry identifiable traces of speakers (tseng2021membership). Related studies further show that models can infer sensitive personal attributes from audio, a capability known as audio private attribute profiling wang2025man. More recent efforts examine interactive vulnerabilities in LALMs, showing that audio-based jailbreak prompts can bypass safety filters and that backdoor triggers can be injected during training to manipulate model behavior (yang2025can; audiojailblogs). To mitigate these threats, several approaches have emerged: representation-level anonymisation obscures speaker identity while retaining linguistic meaning (tran2023privacy); machine unlearning techniques aim to erase specific speakers or utterances post-training (koudounas2025alexa; cheng2025speech); and front-end adversarial defences attempt to detect or filter malicious audio (alexos2025defending).
@@ 3.1 Overview
SH-Bench is a benchmark for evaluating and improving bystander privacy protection in LALMs through the task of selective hearing. It is designed to assess whether LALMs can attend to target speakers while ignoring incidental bystander speech using a test set, which is divided into two partitions: a real-audio partition and a synthetic-audio partition. Beyond evaluation, SH-Bench also supports model improvement by providing a training set for fine-tuning.
@@ 3.2 Test Set Construction
Script Generation. We used GPT-4o (hurst2024gpt) to generate separate scripts for a main speaker and a bystander in each scene. The main speaker’s script consists of structured, purposeful content intended for interaction with a LALM (e.g., podcast monologues, virtual meetings, casual self-talk, or voice assistant queries). In contrast, the bystander’s script contains unrelated, informal, and often sensitive speech (e.g., personal conversations, health disclosures, or travel plans), designed to reflect incidental background speech. Main speaker scripts are longer and more coherent, while bystander scripts consist of a few short turns. The prompt used to generate these scripts is provided in Appendix A.1.
@@
Speaker Recruiting. We recruited English-speaking participants (aged 18+) in pairs through the Prolific  [^3] https://www.prolific.com/ platform to record audio for each scenario. Each pair was assigned the roles of main speaker and bystander. Participants received detailed instructions, and recordings were captured using their personal devices to preserve the acoustic variability of everyday environments. All recordings were manually reviewed by two authors for clarity and consistency with the scripts before inclusion in the dataset. Ethical considerations are detailed in §9.
@@
Generating the Synthetic Scenarios: We first find audio segments of 2-3 minutes long where the speaker is speaking more than 70% of the time based on the rich transcription provided in AMI. These segments are used as the main speaker. We then find segments of 20-50 seconds with dense information content as the bystander audio segments. We attenuate bystander audio by -10dB relative to the main speaker, and mix it into the main segment at a randomly selected point. Meeting room reverberation is added during the mixing process.
@@ 4 Bystander Privacy Fine-Tuning
Specifically, following the same data creation pipeline for the synthetic audio partition (§3.2), we scale up and curate a training set containing 3768 audio mixtures with 75k questions, including both MCQs and open-ended questions, with 37.5k related to the main speaker and 37.5k to the bystander. Each question in the training set has a pair of instructions, where one is to answer questions in general and another is to refuse if the question is about the bystander. As a result, this process will not only encourage the model to learn to distinguish and protect bystander privacy, but also enhance its performance in multi-speaker scenarios in general.
@@ 5.1 Models
Open-source LLMs: We investigate popular LLMs with multi-speaker audio understanding abilities, including Qwen-2.5-Omni 3B and 7B qwen25, Llama-Omni2-14B fang2025llamaomni2, Step-Audio-2-mini stepaudio2 and Kimi-Audio-7B-Instruct kimiaudio. As a screening process, models were instructed to distinguish how many speakers are there in the audio, and the above models were selected that can give the correct number most of the times. We closely follow the instructions provided on the official repositories, including the exact same system prompt and suggested decoding configurations for each model. Inference code is provided at https://github.com/Elocinacademia/SelectiveHearing-Bench.git.
Proprietary LLMs: We study Gemini 2.5 Pro gemini25 and GPT-4o-audio-preview hurst2024gpt as two powerful proprietary models with audio perception abilities.
@@ 5.2 Evaluation
The questions in SH-Bench are 5-way classification tasks with one of the options being “I don’t know" (or equivalent phrases). We evaluate the models under two different modes: (i). General mode, where we instruct the model to answer the question given by the audio, and the model is expected to answer all questions correctly regardless of whether they concern the main speaker or the bystander. (ii). Selective mode where we instruct the model that there is a main speaker in the foreground and provide a content-conditioned description of this speaker (see §3.2). The model is also instructed to only listen to the main speaker and ignore any background speech, so it should choose the IDK option for the questions related to what the bystander said, and should give correct answer for the main speaker. The prompts used in evaluation are provided in Appendix A.1, and the evaluation procedure is shown with examples in Fig. 3.
Figure 3: Illustration of how accuracy is measured for the main speaker and the bystander in two modes. The main speaker must give the correct answer in both modes, whereas the bystander is only correct when the LALM gives the correct answer in general mode but selects “IDK” in the selective mode.
@@ 6.1 Main Benchmark Results
For selective mode, the task is mainly to test if the model can follow instructions and clearly distinguish which content is said by the main speaker and which is by the bystander. Since we provide “I don’t know" as one option, an over-reserved model may always choose “I don’t know" without actually understand the content (e.g. Llama-Omni-2) and often performs worse on main speakers as well under selective mode. Therefore, we report SE to reflects the selective hearing ability of all models by balancing all 4 accuracies. For existing models, Gemini 2.5 Pro achieved the best performance with SE of 75.8%, clearly higher than any other systems.
Real vs. synthetic scenarios: Accuracies on real and synthetic scenarios are given in Table LABEL:tab:real_vs_synth in Appendix A.2. Since real scenarios are more challenging, we observed consistently lower accuracies on real scenarios for bystander speech under general mode, and vice versa under selective mode. The main speaker accuracies have negligible differences under real and synthetic scenarios.
@@
BPFT achieved Consistently Better Performance. After fine-tuning with BPFT, both Step-Audio-2-mini and Qwen-2.5-Omni obtained significant performance improvements on bystander. The model is more capable of understanding the speech content of a background speaker in various environment, and more importantly, it learns to protect bystander privacy when instructed to do so. As a result, Step-Audio-2-mini with BPFT achieved the best SE across the table, with 15.9% absolute SE increase compared to Gemini 2.5 Pro. Remarkably, by just fine-tuning with synthetic scenario data, the model learns to protect bystander privacy at a 96.1% accuracy without influencing the model performance on main speakers.
However, BPFT is by no means a perfect privacy protection mechanism. Although better SE was achieved, we notice that BPFT might cause small degradation to the main speaker accuracy as shown by the Qwen-2.5-Omni results.

### v2

# Protecting Bystander Privacy via Selective Hearing in Audio LLMs

@@ Abstract
Audio Large language models (LLMs) are increasingly deployed in the real world, where they inevitably capture speech from unintended nearby bystanders, raising privacy risks that existing benchmarks and defences did not consider. We introduce SH-Bench, the first benchmark designed to evaluate selective hearing: a model’s ability to attend to an intended main speaker while refusing to process or reveal information about incidental bystander speech. SH-Bench contains 3,968 multi-speaker audio mixtures, including both real-world and synthetic scenarios, paired with 77k multiple-choice questions that probe models under general and selective operating modes. In addition, we propose Selective Efficacy (SE), a novel metric capturing both multi-speaker comprehension and bystander-privacy protection. Our evaluation of state-of-the-art open-source and proprietary LLMs reveals substantial bystander privacy leakage, with strong audio understanding failing to translate into selective protection of bystander privacy. To mitigate this gap, we also present Bystander Privacy Fine-Tuning (BPFT), a novel training pipeline that teaches models to refuse bystander-related queries without degrading main-speaker comprehension. We show that BPFT yields substantial gains, achieving an absolute 47% higher bystander accuracy under selective mode and an absolute 16% higher SE compared to Gemini 2.5 Pro, which is the best audio LLM without BPFT. Together, SH-Bench and BPFT provide the first systematic framework for measuring and improving bystander privacy in audio LLMs. [^1] https://huggingface.co/datasets/BrianatCambridge/SelectiveHearingBench
Protecting Bystander Privacy via Selective Hearing in Audio LLMs
@@ 1 Introduction
Audio Large language models (LLMs), especially the recent efforts including Speech-LLaMA (wu2023decoder), SALMONN (tang2023salmonn), BLSP (wang2023blsp), and Qwen-Audio (chu2023qwen), extend the capabilities of text-based LLMs to the acoustic domain. As audio LLMs are deployed in real-world settings such as voice assistants and wearable devices metaglasses; geminilive, they passively capture open-domain speech in uncontrolled environments, which inevitably introduces significant privacy risks. Human voices contain sensitive acoustic attributes such as timbre, pitch, and prosody that can reveal identity, emotional state, and health conditions (nautsch2019gdpr; backstrom2023privacy; wang2025man; aloufi2021configurable). When trained on large-scale real-world speech corpora, audio LLMs often inadvertently memorise this information hartmann2023sok; mccoy2023much, leading to potential exposure or re-identification chen2023slmia. Moreover, prior work further shows that LLMs are vulnerable to various privacy attacks  (tseng2021membership; carlini2021extracting; yang2025can; audiojailblogs) which amplify the risks of sensitive information leakage.
@@
Figure 1: An illustration of the bystander privacy challenge in audio LLMs. The primary speaker interacts with an audio LLM, while nearby bystanders may be unintentionally recorded and could unknowingly reveal private information. To protect bystander privacy, the audio LLM should attend only to the primary speaker and refuse to answer any queries concerning bystanders.
@@
Yet, existing mitigation efforts focus primarily on active users (tran2023privacy; koudounas2025alexa; cheng2025speech; alexos2025defending), those who knowingly interact with the model. In contrast, a significant and overlooked group in real-world deployment contexts are bystanders: individuals whose speech is incidentally captured without their knowledge, consent, or intent to engage [^2] In social science, bystander typically refers to an observer of an event without active participation (e.g, “bystander effect” in zapata2024bystanders). Here we use the definition and research focus established in smart-device contexts yao2019privacy; saqib2025bystander, viewed through a privacy lens that highlights risks when data is passively captured or shared.. Bystanders face the same privacy risks as active users, but they neither control nor even know how their speech is processed. This disconnect raises a critical question: Can audio LLMs be designed to selectively attend to intended input while refusing to expose bystander information? As illustrated in Fig.1, the audio LLM should refuse any request targeting a bystander who may unknowingly disclose private speech.
@@
To address the critical gap of bystander privacy in audio LLMs, that is the quantification, comparison, or evaluation of bystander privacy in audio LLMs, this paper proposes SH-Bench, a Selective Hearing Benchmark. SH-Bench is the first benchmark for assessing the capability of audio LLMs to protect bystander privacy through selective hearing. SH-Bench consists of multi-speaker audio samples with five-way classification tasks where one of the option designed to be “I don’t know”. In addition, an evaluation framework is designed for SH-Bench which allows the assessment of both model comprehension abilities in multi-speaker scenarios and bystander privacy protection, with a unified criterion: Selective Efficacy (SE), a novel metric that we propose. Moreover, we introduce the Bystander Privacy Fine-Tuning (BPFT), providing training data intended to enhance bystander protection in audio LLMs. Overall, our results reveal a substantial lack of bystander privacy protection in existing audio LLMs without fine-tuning. The main contributions of this paper are summarised as follows.
@@
  • We propose SH-Bench, the first selective hearing benchmark for audio LLMs to assess bystander privacy protection in these LLMs in multi-speaker environments.
@@
  • We contribute the evaluation framework for SH-Bench, including two different operation modes and speaking parties. We also propose SE as a unified metric balancing model comprehension abilities and bystander privacy protection.
@@
  • We propose a bystander privacy fine-tuning (BPFT) pipeline from training data curation to supervised fine-tuning to enhance bystander privacy in audio LLMs. BPFT achieves substantial improvements on bystander privacy protection, with an absolute 47% higher bystander accuracy under selective mode and an absolute 15.9% higher SE compared to Gemini 2.5 Pro.
@@ 2.1 Multi-Speaker Benchmarks
It is common for speech benchmarks to be either fully multi-speaker or to include substantial multi-speaker segments, as speaker diarization is one of the core representative tasks in speech processing. We categorise existing multi-speaker benchmarks into two groups: (i) benchmarks primarily designed for non-privacy related tasks such as automatic speech recognition (ASR), spoken question answering, speaker diarisation and speech separation, including traditional benchmarks  kraaij2005ami; cosentino2020librimix; godfrey1992switchboard; zeinali2018deepmine; garcia2019speaker, more recent benchmarks tailored to audio LLMs huang2024dynamic; sakshi2024mmau; yue2024mmmu; sun2025audioset; wang2025msu, and audio-visual multi-speaker benchmarks yang2025audio; tseng2024av to evaluate tasks that require joint audio-visual understanding; and (ii) safety- and privacy-oriented audio benchmarks focus on model robustness, safety risks, and privacy leakage in multi-speaker settings. For instance, the multi-speaker anonymisation benchmark in miao2025benchmark examines privacy risks and mitigation strategies in overlapping conversational audio, and SACRED-Bench yang2025speech, the first multi-speaker jailbreak benchmark, features dialogues where harmful instructions are embedded within or alongside benign speech. However, these benchmarks focus solely on active speakers and overlook bystanders, leaving a critical gap that we address in this study.
2.2 Privacy Risks with Audio LLMs
Privacy research on audio LLMs has largely focused on protecting the primary speaker, without distinguishing intended users from bystanders. Early work reveals that both end-to-end ASR and self-supervised speech encoders can leak training membership information through black-box queries, demonstrating that speech representations themselves carry identifiable traces of speakers (tseng2021membership). Related studies further show that audio models can infer sensitive personal attributes, known as audio private attribute profiling wang2025man, and exposes interactive vulnerabilities such as audio-based jailbreaks and training-time backdoor triggers (yang2025can; audiojailblogs). To mitigate these risks, prior approaches include representation-level anonymisation  (tran2023privacy), machine unlearning (koudounas2025alexa; cheng2025speech), and front-end adversarial defences  (alexos2025defending). However, to the best of our knowledge, no prior work has tackled the problem of bystander privacy in audio LLMs.
@@ 3.1 Overview
SH-Bench is a benchmark that contains both a test set and a training set, enabling the evaluation and improvement of bystander privacy protection in audio LLMs through the task of selective hearing. It is designed to assess whether audio LLMs can attend only to target speakers using a test set, which is divided into two partitions: a real-audio partition and a synthetic-audio partition. Beyond evaluation, SH-Bench also supports model improvement by providing a separate training set for fine-tuning (its construction is described in §4).
@@ 3.2 Test Set Construction
Script Generation. We used GPT-4o (hurst2024gpt) to generate separate scripts for a main speaker and a bystander in each scene. The main speaker’s script consists of structured, purposeful content intended for interaction with an audio LLM (e.g., podcast monologues, virtual meetings, casual self-talk, or voice assistant queries). In contrast, the bystander’s script contains unrelated, informal, and often sensitive speech (e.g., personal conversations, health disclosures, or travel plans), designed to reflect incidental background speech. Main speaker scripts are longer and more coherent, while bystander scripts consist of a few short turns. The prompt used to generate these scripts is provided in Appendix A.1.
@@
Speaker Recruiting. We recruited English-speaking participants (18+) in pairs viaProlific [^3] https://www.prolific.com/ to record audio for each scenario. Each pair was assigned main speaker and bystander roles, followed detailed instructions, and recorded using personal devices to capture real-world acoustic variability. All recordings were manually reviewed by two authors for clarity and consistency with the scripts before inclusion in the dataset. Ethical considerations are detailed in §9.
@@
Generating the Synthetic Scenarios: We first find audio segments of 2-3 minutes long where the speaker is speaking more than 70% of the time based on the rich transcription provided in AMI. These segments are used as the main speaker. We then find segments of 20-50 seconds with dense information content as the bystander audio segments. We attenuate bystander audio by -10dB relative to the main speaker, reflecting practical scenarios where bystanders are typically farther from the microphone, and mix it into the main segment at a randomly selected point. Meeting room reverberation is added during the mixing process.
@@ 4 Bystander Privacy Fine-Tuning
Specifically, following the same data creation pipeline for the synthetic audio partition (§3.2), we scale up and curate a training set containing 3,768 audio mixtures with 75k questions, including both MCQs and open-ended questions, with 37.5k related to the main speaker and 37.5k to the bystander. Each question in the training set has a pair of instructions, where one is to answer questions in general and another is to refuse if the question is about the bystander. As a result, this process will not only encourage the model to learn to distinguish and protect bystander privacy, but also enhance its performance in multi-speaker scenarios in general.
@@ 5.1 Models
Open-source audio LLMs: We investigate popular LLMs with multi-speaker audio understanding abilities, including Qwen-2.5-Omni 3B and 7B qwen25, Llama-Omni2-14B fang2025llamaomni2, Step-Audio-2-mini stepaudio2 and Kimi-Audio-7B-Instruct kimiaudio. As a screening process, models were instructed to distinguish how many speakers are there in the audio, and the above models were selected that can give the correct number most of the times. We closely follow the instructions provided on the official repositories, including the exact same system prompt and suggested decoding configurations for each model. Inference code is provided at https://github.com/Elocinacademia/SelectiveHearing-Bench.git.
Proprietary audio LLMs: We study Gemini 2.5 Pro gemini25 and GPT-4o-audio-preview hurst2024gpt as two powerful proprietary models with audio perception abilities.
@@ 5.2 Evaluation
SH-Bench evaluation questions use a 5-way classification format, including an “I don’t know” option (or equivalent). With these questions, we evaluate the models under two different modes: i) General mode, where we instruct the model to answer the question given by the audio, and the model is expected to answer all questions correctly regardless of whether they concern the main speaker or the bystander; and ii) Selective mode, where we instruct the model that there is a main speaker in the foreground and provide a content-conditioned description of this speaker (see §3.2). The model is also instructed to only listen to the main speaker and ignore any background speech, so it should choose the IDK option for the questions related to what the bystander said, and should give correct answer for the main speaker. The prompts used in the evaluation are provided in Appendix A.1, and the evaluation procedure is shown with examples in Fig. 3.
Figure 3: Illustration of how accuracy is measured for the main speaker and the bystander in two modes. The main speaker must give the correct answer in both modes, whereas the bystander is only correct when the audio LLM gives the correct answer in general mode but selects “IDK” in the selective mode.
@@ 6.1 Main Benchmark Results
For selective mode, the task is mainly to test if the model can follow instructions and clearly distinguish which content is said by the main speaker and which is by the bystander. Since we provide “I don’t know" as one option, an over-reserved model may always choose “I don’t know" without actually understand the content (e.g. Llama-Omni-2) and often performs worse on main speakers as well under selective mode. Therefore, we report SE to reflect the selective hearing ability of all models by balancing all 4 accuracies. For existing models, Gemini 2.5 Pro achieved the best performance with SE of 75.8%, clearly higher than any other systems.
Real vs. synthetic scenarios: Accuracies on real and synthetic scenarios are given in Table 6 in Appendix A.2. Since real scenarios are more challenging, we observed consistently lower accuracies on real scenarios for bystander speech under general mode, and vice versa under selective mode. The main speaker accuracies have negligible differences under real and synthetic scenarios.
@@
BPFT achieved Consistently Better Performance. After fine-tuning with BPFT, both Step-Audio-2-mini and Qwen-2.5-Omni obtained significant performance improvements on bystander. The model is more capable of understanding the speech content of a background speaker across environments, and more importantly, it learns to protect bystander privacy when instructed to do so. As a result, Step-Audio-2-mini with BPFT achieved the best SE across the table, with an absolute 15.9% higher SE compared to Gemini 2.5 Pro (the best one without BPFT). Remarkably, by just fine-tuning with synthetic scenario data, the model learns to protect bystander privacy at a 96.1% accuracy without influencing the model performance on main speakers.
However, BPFT is by no means a perfect privacy protection mechanism. Although better SE was achieved, we notice that BPFT might cause a slight degradation to the main speaker accuracy as shown by the Qwen-2.5-Omni results. This is, however, not observed in Step-Audio-2-mini where main speaker accuracy actually improves slightly. Importantly, in both cases the increase in bystander privacy is very substantial, of circa between 50 to 60 percentage points.